OpenStudy (anonymous):

What mass of silver would be deposited if a 6.5 A current is passed through an aqueous solution of silver nitrate for 8.0 minutes?

OpenStudy (anonymous):

@PANTIEPOWER

OpenStudy (anonymous):

0.0 dunno know lol sorry @abb0t Mr. king

OpenStudy (anonymous):

@abb0t anything?

OpenStudy (abb0t):

Well, FIRST, you need to set up your \(\sf \color{red}{balanced~ chemical ~equation}\). Then, you need to find the amount of moles, divide by faradays constant. which is \(\sf 9.65 \times 10^4 ~\frac{C}{mol}\) Then multiply, which eliminates the Ampere units giving you moles of electron.

OpenStudy (abb0t):

You need to get it in seconds also, by the way.

OpenStudy (abb0t):

convert 8.0 mins \(\sf \rightarrow\) seconds Then, once you have your moles of electron, use dimensional analysis to find your moles of silver.

OpenStudy (anonymous):

I'm really confused haha @abb0t

OpenStudy (abb0t):

let's go step by step, start by writing down the balanced chemical equation, can you do that?

OpenStudy (anonymous):

Cu + AgNO3 --> CuNO3 + Ag ?

OpenStudy (anonymous):

wait no Cu + 2AGNo3 --> Cu(No3)2 + 2AG

OpenStudy (anonymous):

@abb0t @aaronq

OpenStudy (abb0t):

Good, now convert minutes to seconds and divide! This will give you coulombs, which you can then use to divide by \(faraday's~constant\), which i posted up above, the 96500..

OpenStudy (abb0t):

This will give you moles. Then, using your balanced chemical equation, convert from what you have to what you want. Do you follow me still?

OpenStudy (anonymous):

so 8 mins = 480 seconds but what am I dividing?

OpenStudy (anonymous):

@abb0t

OpenStudy (abb0t):

Follow the units, it will tell you what you need to divide by.

OpenStudy (abb0t):

I believe it should be Ampere per seconds.

OpenStudy (anonymous):

so 6.5/480 = 0.013542 then that divided by F constant is 1.40331606e-7 what next?

OpenStudy (anonymous):

@abb0t

OpenStudy (anonymous):

@Abmon98 ?

OpenStudy (abb0t):

yes.

OpenStudy (anonymous):

what's next? @abb0t

OpenStudy (abb0t):

Did you divide by faradays constant?

OpenStudy (anonymous):

yes

OpenStudy (abb0t):

notice that you are left with moles.

OpenStudy (anonymous):

1.40331606e-7 moles?

OpenStudy (abb0t):

No, not that one. I already provded it for you up above!

OpenStudy (anonymous):

I'm so sorry I'm so lost hahaha @abb0t

OpenStudy (abb0t):

Look up above. 96500...

OpenStudy (anonymous):

right the 9.65×10^4

OpenStudy (anonymous):

I divided 6.5/480 = 0.013542 and then I divided that by faraday and i got 1.40331606e-7

OpenStudy (abb0t):

\(\large{\frac{0.013542}{96500}}\) = mol \(\sf \times \frac{mol~want}{mol~have}\)

OpenStudy (abb0t):

This is where you will utilize your balanced chemical equation to convert! Do you follow me?

OpenStudy (anonymous):

ok yes I follow you now

OpenStudy (abb0t):

Now that you have your moles, convert to moles of Ag, since you have Ag(NO3), note that it is a 2:2 ratio (i think if you wrote it correctly - make sure it is). so you have \(\sf \frac{2~mol~ Ag}{2~mol~Ag(NO3)}\) = mol Ag now, you can use molar mass to convert and find the mass. Remember that molar mass is grams per mol, so make sure that you look at your periodic table, and look for Ag - silver.

OpenStudy (anonymous):

so \[\frac{ 0.013542 }{ 96500 } = mol Ag\] and then \[1.40331606\times10^{-7} \times 107.86820\]

OpenStudy (anonymous):

@abb0t

OpenStudy (anonymous):

I got a wrong answer... hmm I got 0.00001513731 but the answer choices are as follows:

OpenStudy (abb0t):

I got one of those answers.

OpenStudy (abb0t):

\(\large{\frac{6.5 \times 480~s}{96500 }}\) = __________ mol your mole is incorrect. I don't know where you're getting such a low mole value!

OpenStudy (abb0t):

check your math again.

OpenStudy (abb0t):

You should get 3.459...

OpenStudy (anonymous):

@abb0t this helped me a lot